\begin{thebibliography}{10}

\bibitem{criu}
{CRIU, a project to implement checkpoint/restore functionality for Linux.}
\newblock \url{https://criu.org}.
\newblock [Online].

\bibitem{dmtcp_faq_2021}
{DMTCP (FAQ): Distributed MultiThreaded CheckPointing}.
\newblock \url{http://dmtcp.sourceforge.net/FAQ.html}.
\newblock [Retrieved January 18, 2021. Online].

\bibitem{moab_fairshare2019}
{Moab Adaptive Computing Suite Administrator’s Guide - v. 5.4}.
\newblock \url{http://docs.adaptivecomputing.com/macs/6.3fairshare.php}.
\newblock [Retrieved January 18, 2021. Online].

\bibitem{negevhpc}
{NegevHPC Project}.
\newblock \url{www.negevhpc.com}.
\newblock [Online].

\bibitem{pyslurm}
{PySlurm: Slurm Interface for Python}.
\newblock \url{https://github.com/PySlurm/pyslurm}.
\newblock [Online].

\bibitem{slurm}
{SLURM}.
\newblock \url{https://slurm.schedmd.com}.
\newblock [Online].

\bibitem{hoopes_slurm_2019}
Slurm {Priority}, {Fairshare} and {Fair} {Tree}.
\newblock
  \url{https://slurm.schedmd.com/SLUG19/Priority\_and\_Fair\_Trees.pdf}.
\newblock [Online].

\bibitem{slurm_classic_fairshare}
Slurm's classic fairshare algorithm.
\newblock \url{slurm.schedmd.com/classic\_fair\_share.html}.
\newblock [Online].

\bibitem{top500}
{TOP500 List, June 2020}.
\newblock \url{www.top500.org/lists/top500/2020/06/}.
\newblock [Online].

\bibitem{raas}
Orna Agmon Ben-Yehuda, Muli Ben-Yehuda, Assaf Schuster, and Dan Tsafrir.
\newblock The rise of raas: The resource-as-a-service cloud.
\newblock {\em Commun. ACM}, 57(7):76–84, July 2014.

\bibitem{anderson2019assise}
Thomas~E Anderson, Marco Canini, Jongyul Kim, Dejan Kosti{\'c}, Youngjin Kwon,
  Simon Peter, Waleed Reda, Henry~N Schuh, and Emmett Witchel.
\newblock Assise: Performance and availability via nvm colocation in a
  distributed file system.
\newblock {\em arXiv preprint arXiv:1910.05106}, 2019.

\bibitem{assise2020}
Thomas~E. Anderson, Marco Canini, Jongyul Kim, Dejan Kosti{\'c}, Youngjin Kwon,
  Simon Peter, Waleed Reda, Henry~N. Schuh, and Emmett Witchel.
\newblock Assise: Performance and availability via client-local {NVM} in a
  distributed file system.
\newblock In {\em 14th {USENIX} Symposium on Operating Systems Design and
  Implementation ({OSDI} 20)}, pages 1011--1027. {USENIX} Association, November
  2020.

\bibitem{ansel2009dmtcp}
Jason Ansel, Kapil Arya, and Gene Cooperman.
\newblock {DMTCP}: Transparent checkpointing for cluster computations and the
  desktop.
\newblock In {\em 2009 IEEE International Symposium on Parallel \& Distributed
  Processing (IPDPS'09)}, pages 1--12, Rome, Italy, 2009. IEEE.

\bibitem{breuer2019petaflop}
Alexander Breuer, Yifeng Cui, and Alexander Heinecke.
\newblock Petaflop seismic simulations in the public cloud.
\newblock In {\em International Conference on High Performance Computing},
  pages 167--185. Springer, 2019.

\bibitem{buyya2018manifesto}
Rajkumar Buyya, Satish~Narayana Srirama, Giuliano Casale, Rodrigo Calheiros,
  Yogesh Simmhan, Blesson Varghese, Erol Gelenbe, Bahman Javadi, Luis~Miguel
  Vaquero, Marco~AS Netto, et~al.
\newblock A manifesto for future generation cloud computing: Research
  directions for the next decade.
\newblock {\em ACM computing surveys (CSUR)}, 51(5):1--38, 2018.

\bibitem{cao_transparent_2013}
Jiajun Cao, Kapil Arya, and Gene Cooperman.
\newblock Transparent {Checkpoint}-{Restart} over {InfiniBand}.
\newblock {\em HPDC 2014 - Proceedings of the 23rd International Symposium on
  High-Performance Parallel and Distributed Computing}, December 2013.

\bibitem{cui2020survey}
Delong Cui, Zhiping Peng, Qirui Li, Jieguang He, Lizi Zheng, and Yiheng Yuan.
\newblock A survey on cloud workflow collaborative adaptive scheduling.
\newblock In {\em Advances in Computer, Communication and Computational
  Sciences}, pages 121--129. Springer, 2020.

\bibitem{dolev_no_2011}
Danny Dolev, Dror Feitelson, Joseph Halpern, Raz Kupferman, and Nati Linial.
\newblock No {Justified} {Complaints}: {On} {Fair} {Sharing} of {Multiple}
  {Resources}.
\newblock {\em ITCS 2012 - Innovations in Theoretical Computer Science
  Conference}, June 2011.

\bibitem{mualem1998}
D.~G. {Feitelson} and A.~M. {Weil}.
\newblock Utilization and predictability in scheduling the ibm sp2 with
  backfilling.
\newblock In {\em Proceedings of the First Merged International Parallel
  Processing Symposium and Symposium on Parallel and Distributed Processing},
  pages 542--546, 1998.

\bibitem{ferretti2019cloud}
Marco Ferretti and Luigi Santangelo.
\newblock {\em Cloud vs On-Premise HPC: a model for comprehensive cost
  assessment}.
\newblock press, 2019.

\bibitem{hargrove2006berkeley}
Paul~H Hargrove and Jason~C Duell.
\newblock Berkeley lab checkpoint/restart (blcr) for linux clusters.
\newblock In {\em Journal of Physics: Conference Series}, volume~46, page 494,
  2006.

\bibitem{henry_unix_1984}
G.~J. Henry.
\newblock The {UNIX} {System}: {The} {Fair} {Share} {Scheduler}.
\newblock {\em AT\&T Bell Laboratories Technical Journal}, 63(8):1845--1857,
  1984.
\newblock \_eprint:
  https://onlinelibrary.wiley.com/doi/pdf/10.1002/j.1538-7305.1984.tb00068.x.

\bibitem{izraelevitz2019basic}
Joseph Izraelevitz, Jian Yang, Lu~Zhang, Juno Kim, Xiao Liu, Amirsaman
  Memaripour, Yun~Joon Soh, Zixuan Wang, Yi~Xu, Subramanya~R Dulloor, et~al.
\newblock Basic performance measurements of the intel optane dc persistent
  memory module.
\newblock {\em arXiv preprint arXiv:1903.05714}, 2019.

\bibitem{splitfs2019}
Rohan Kadekodi, Se~Kwon Lee, Sanidhya Kashyap, Taesoo Kim, Aasheesh Kolli, and
  Vijay Chidambaram.
\newblock Splitfs: Reducing software overhead in file systems for persistent
  memory.
\newblock In {\em Proceedings of the 27th ACM Symposium on Operating Systems
  Principles}, SOSP '19, page 494–508, New York, NY, USA, 2019. Association
  for Computing Machinery.

\bibitem{kyaw2020scheduling}
Lett~Yi Kyaw and Sabai Phyu.
\newblock Scheduling methods in hpc system.
\newblock In {\em 2020 IEEE Conference on Computer Applications (ICCA)}, pages
  1--6. IEEE, 2020.

\bibitem{lee2004user}
Cynthia~Bailey Lee, Yael Schwartzman, Jennifer Hardy, and Allan Snavely.
\newblock Are user runtime estimates inherently inaccurate?
\newblock In {\em Workshop on Job Scheduling Strategies for Parallel
  Processing}, pages 253--263. Springer, 2004.

\bibitem{liu2020survey}
Haikun Liu, Di~Chen, Hai Jin, Xiaofei Liao, Bingsheng He, Kan Hu, and Yu~Zhang.
\newblock A survey of non-volatile main memory technologies: State-of-the-arts,
  practices, and future directions.
\newblock {\em arXiv preprint arXiv:2010.04406}, 2020.

\bibitem{meuer2014top500}
Hans~Werner Meuer, Erich Strohmaier, Jack Dongarra, and Horst~D Simon.
\newblock The top500: History, trends, and future directions in high
  performance computing.
\newblock 2014.

\bibitem{netto2018hpc}
Marco~AS Netto, Rodrigo~N Calheiros, Eduardo~R Rodrigues, Renato~LF Cunha, and
  Rajkumar Buyya.
\newblock Hpc cloud for scientific and business applications: taxonomy, vision,
  and research challenges.
\newblock {\em ACM Computing Surveys (CSUR)}, 51(1):1--29, 2018.

\bibitem{Niu2012EmployingCT}
Shuangcheng Niu, Jidong Zhai, Xiaosong Ma, Mingliang Liu, Y.~Zhai, W.~Chen, and
  W.~Zheng.
\newblock Employing checkpoint to improve job scheduling in large-scale
  systems.
\newblock In {\em JSSPP}, 2012.

\bibitem{rodriguez2019job}
Manuel Rodr{\'\i}guez-Pascual, Jiajun Cao, Jos{\'e}~A Mor{\'\i}{\~n}igo, Gene
  Cooperman, and Rafael Mayo-Garc{\'\i}a.
\newblock Job migration in hpc clusters by means of checkpoint/restart.
\newblock {\em The Journal of Supercomputing}, 75(10):6517--6541, 2019.

\bibitem{sadooghi2015understanding}
Iman Sadooghi, Jes{\'u}s~Hern{\'a}ndez Martin, Tonglin Li, Kevin Brandstatter,
  Ketan Maheshwari, Tiago Pais~Pitta de~Lacerda~Ruivo, Gabriele Garzoglio,
  Steven Timm, Yong Zhao, and Ioan Raicu.
\newblock Understanding the performance and potential of cloud computing for
  scientific applications.
\newblock {\em IEEE Transactions on Cloud Computing}, 5(2):358--371, 2015.

\bibitem{scargall2020introducing}
Steve Scargall.
\newblock Introducing the persistent memory development kit.
\newblock In {\em Programming Persistent Memory}, pages 63--72. Springer, 2020.

\bibitem{steffen2019better}
Craig~P Steffen.
\newblock A better way of scheduling jobs on hpc systems: Simultaneous
  fair-share.
\newblock 2019.

\bibitem{uruchi2020game}
Wilmer Uruchi~Ticona.
\newblock Game theoretic analysis of the slurm scheduler model.
\newblock 2020.

\bibitem{nova2017}
Jian Xu, Lu~Zhang, Amirsaman Memaripour, Akshatha Gangadharaiah, Amit Borase,
  Tamires~Brito Da~Silva, Steven Swanson, and Andy Rudoff.
\newblock Nova-fortis: A fault-tolerant non-volatile main memory file system.
\newblock In {\em Proceedings of the 26th Symposium on Operating Systems
  Principles}, SOSP '17, page 478–496, New York, NY, USA, 2017. Association
  for Computing Machinery.

\bibitem{yang2019orion}
Jian Yang, Joseph Izraelevitz, and Steven Swanson.
\newblock Orion: A distributed file system for non-volatile main memory and
  rdma-capable networks.
\newblock In {\em 17th $USENIX$ Conference on File and Storage Technologies
  ($FAST$ 19)}, pages 221--234, 2019.

\bibitem{yoo2003slurm}
Andy~B Yoo, Morris~A Jette, and Mark Grondona.
\newblock Slurm: Simple linux utility for resource management.
\newblock In {\em Workshop on Job Scheduling Strategies for Parallel
  Processing}, pages 44--60. Springer, 2003.

\bibitem{yuan_pv-easy_2010}
Yulai Yuan, Guangwen Yang, Yongwei Wu, and Weimin Zheng.
\newblock {PV}-{EASY}: a strict fairness guaranteed and prediction enabled
  scheduler in parallel job scheduling.
\newblock In {\em Proceedings of the 19th {ACM} {International} {Symposium} on
  {High} {Performance} {Distributed} {Computing}}, {HPDC} '10, pages 240--251,
  New York, NY, USA, June 2010. Association for Computing Machinery.

\end{thebibliography}

\documentclass[letterpaper, 10 pt, conference]{ieeeconf} 
\IEEEoverridecommandlockouts                              
\usepackage{lipsum}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{algorithm}
\usepackage{url}

\usepackage{lipsum}% http://ctan.org/pkg/lipsum

\usepackage[noend]{algpseudocode}
\usepackage{algpseudocode}
\usepackage[bottom]{footmisc}
\usepackage{tikz}

\usetikzlibrary{shapes.arrows, fadings}
\usetikzlibrary{decorations.text}

\usepackage{cite}

\usepackage{mathtools}

\DeclarePairedDelimiter\floor{\lfloor}{\rfloor}
\usepackage[misc]{ifsym}
\newcommand{\corrAuthor}{$^{\textrm{\Letter}}$}
\newcommand{\MyComment}[1]{\hspace*{\fill}\Comment{\parbox[t]{.4\linewidth}{#1}}}

\usepackage{flushend}

\begin{document}

\title{\LARGE \bf
Optimized Memoryless Fair-Share HPC Resources Scheduling \\ using Transparent Checkpoint-Restart Preemption
}

\author{
    Kfir Zvi$^{1}$, Gal Oren$^{2}$\corrAuthor
    \thanks{
        $^{1}$K. Zvi is with the Department of Computer Science, Ben-Gurion University of the Negev, Be'er-Sheva, Israel
        \& Israel Atomic Energy Commission, P.O.B. 7061, Tel Aviv, Israel
        \newline
        {\tt\small  zvikf@post.bgu.ac.il}
    }
    \thanks{
        $^{2}$G. Oren is with the Department of Computer Science, Technion – Israel Institute of Technology, Haifa 32000, Israel
        \& the Scientific Computing Center, Nuclear Research Center-Negev, P.O.B. 9001, Be'er-Sheva, Israel
        \newline
        {\tt\small  galoren@technion.ac.il}
    }

    \thanks{
        \newline
        * This work was supported by Pazy grant 226/20. Computational support was provided by the NegevHPC project~\cite{negevhpc}. The authors would like to thank Matan Rusanovsky, Harel Levin, Hagit Attiya and Danny Hendler for their fruitful comments.
        \newline
    }
}

\maketitle

\begin{abstract}
Common resource management methods in supercomputing systems usually include hard divisions, capping, and quota allotment. Those methods, despite their ‘advantages’, have some known serious disadvantages including unoptimized utilization of an expensive facility, and occasionally there is still a need to dynamically reschedule and reallocate the resources. Consequently, those methods involve bad supply-and-demand management rather than a free market playground that will eventually increase system utilization and productivity. In this work, we propose the newly Optimized Memoryless Fair-Share HPC Resources Scheduling using Transparent Checkpoint-Restart Preemption, in which the social welfare increases using a free-of-cost interchangeable proprietary possession scheme. Accordingly, we permanently keep the status-quo in regard to the fairness of the resources distribution while maximizing the ability of all users to achieve more CPUs and CPU hours for longer period without any non-straightforward costs, penalties or additional human intervention.    
\end{abstract}

\section{Introduction}
Demand for High-Performance Computing (HPC) resources has been skyrocketing over the last decade~\cite{meuer2014top500}. Consequently, said resources became more available over the cloud, affordable (due to competition~\cite{raas}), and immediately supplied~\cite{buyya2018manifesto,netto2018hpc,sadooghi2015understanding}.
Nevertheless, HPC on the cloud is limited~\cite{ferretti2019cloud, breuer2019petaflop}: (1) Cloud hardware is general purposed, and not tailor-made like a supercomputer. Hence, it is insufficient for large-scale coupled computations; (2) The current commercial tariff paradigm makes it more expensive to run continuous computations on the cloud, compared with on-premise resources. Thus, many R\&D institutions still establish in-house supercomputing facilities~\cite{top500}. However, unlike cloud resources, which are dynamically assigned according to an economic model~\cite{cui2020survey}, supercomputing facilities are usually assigned by hard divisions or by utilization capping.
These models reflect entitlements of \textit{entities} (groups or individual users)~\cite{kyaw2020scheduling},
that originate in centrally-managed work-plans. An entity expects immediate access to its resources, 
similarly to reserved cloud instances,
but lacking an economic mechanism -- which would have regulated user requests -- congestion is formed. To prevent group starvation, the congestion is usually mitigated by capping usage, static allocation and rationing CPU hours. 
Jobs are prioritized in queues, and low priority jobs may starve, in full accordance with the users' expectations.
These methods suffer from under-utilization of the expensive, divided facility, and from a frequent need 
to manually adjust the entitlements~\cite{uruchi2020game}. Therefore, there is a need for a scheduling mechanism that will alleviate  the barriers that prevent full resource pooling, and yet maintain the entities' expectation of \textit{fairness}:
Each entity gets at least its entitlement of CPUs when it has the suitable workload for it~\cite{dolev_no_2011}.

\section{The Optimized Memoryless Fair-Share}

The scheduling and resource allocation is usually performed using schedulers such as SLURM~\cite{slurm,yoo2003slurm,top500}.

A recent development of the SLURM scheduler is the seamless incorporation of the Distributed MultiThreaded Checkpointing (DMTCP)~\cite{ansel2009dmtcp} library, enabling it to transparently Checkpoint and Restart (C/R) a single-host, 
parallel or distributed computation. It does so in user-space, with no modifications to user code or the 
operating system, supporting a variety of HPC languages and infrastructures, including MPI and OpenMP~\cite{rodriguez2019job}.

Following this recent capability, we propose the Optimized Memoryless Fair-Share HPC Resource Scheduling (OMFS) method by enabling transparent resource pooling. 
Each entity is still guaranteed its \textit{CPU entitlement}, while not being limited to certain compute nodes via static allocation, and without computation capping. Instead, underutilized resources are assigned to entities with active demand. \textit{Non-preemptible} jobs can only be run up to the entity's entitlement. To utilize more resources,
jobs must be either \textit{preemptable} (may be killed) or \textit{checkpointable} (can C/R using DMTCP or otherwise). Hence, if an entity does not utilize its full entitlement, it can always increase its utilization up to its entitlement by evicting jobs of entities utilizing more than their allotment. If these jobs are checkpointable, they are transparently checkpointed prior to the eviction, and later restarted. 

\begin{algorithm*}[ht]
\begin{algorithmic}[1]
\normalsize
\Procedure{System Init}{$Users$, $Jobs_{Submitted}$, $Jobs_{Running}$}
\MyComment{Performed once at startup}
\label{sys_init_start}
\State $CPU_{Total} \gets N$ \label{cpu_total_def} \MyComment{Amount of CPUs in the entire system}

\State $CPU_{Idle} \gets 0$ \MyComment{Amount of idle CPUs in the entire system}

\State $Users \gets Users$ \MyComment{Predefined group of system users}

\State $Jobs_{Submitted} \gets Jobs_{Submitted}$ \label{jobs_queues_start} \MyComment{A predefined priority queue}
\State $Jobs_{Running} \gets Jobs_{Running}$ \label{jobs_queues_end} \MyComment{A predefined priority queue}
\State \textbf{\textit{foreach}} user $u \in Users$:
\State \hspace{0.5cm} $u.percent \gets p : p \in [0,100]$ \MyComment{Percentage of CPU allocation per user}
\State \textbf{assert} $\sum_{u \in Users}(u.percent) \leq 100$ \MyComment{Upper bound of allocations' percentages} \label{sys_init_end}
\EndProcedure

\item[]
\Procedure{Job Init}{User u, Job j} \MyComment{Performed for every job creation} \label{job_init_start}

\State $j.priority \gets pr : pr \in N_{0}$ \MyComment{Priority among the jobs of the user only}
\State $j.CPU_{Count} \gets cpus : cpus \in \{x \in N_{0}:0\le x\le N\}$ \MyComment{Amount of requested job's CPUs}
\State $j.user \gets u$ \label{job_init_end}
\EndProcedure

\item[]
\Procedure{Memoryless Fair-Share Scheduler()}{} \label{scheduler_start}
\State \textbf{\textit{while}} True: \MyComment{Continuously try to run jobs from $Jobs_{Submitted}$}
\State \hspace{0.5cm} $J \gets Jobs_{Submitted}.dequeue()$ 
\State \hspace{0.5cm} \textsc{Memoryless Fair-Share Runner}(J) \label{scheduler_end}

\EndProcedure

\item[]
\Procedure{Memoryless Fair-Share Runner}{\textit{Job J}} \label{runner_start}

\State $User_{PAbleJobsCPUCount} \gets \sum_{j \in \{ j \in Jobs_{Running} : j.user = J.user \textit{ \textbf{and} j \textbf{is} preemptable}\}} (j.CPU_{Count})$ \newline \MyComment{Amount of CPUs occupied by user's preemptable jobs}
\State $User_{NonPAbleJobsCPUCount} \gets \sum_{j \in \{ j \in Jobs_{Running} : j.user = J.user \text{\textit{ \textbf{and} j \textbf{is not} preemptable}\}}} (j.CPU_{Count})$ \newline \MyComment{Amount of CPUs occupied by user's non-preemptable jobs}
\State $User_{TotalJobsCPUCount} \gets User_{PAbleJobsCPUCount} + User_{NonPAbleJobsCPUCount}$
\State $User_{EntitledCPUCount} \gets \floor{(J.user.percent / 100)\cdot CPU_{Total}}$

\State \textbf{\textit{if}} $J$ \textit{\textbf{is not} preemptable \textbf{and}} $User_{NonPAbleJobsCPUCount} + J.CPU_{Count} \ge User_{EntitledCPUCount}$: \newline \MyComment{If the non-preemptable jobs exceed the user's entitlement} \newline
\State \hspace{0.5cm} $Jobs_{Submitted}.enqueue(J)$ \MyComment{Do not allow to run}
\State \hspace{0.5cm} \textbf{\textit{return}}
\State \textbf{\textit{elif}} $CPU_{Idle} > J.CPU_{Count}$: \MyComment{If enough resources are available}
\State \hspace{0.5cm} \textbf{\textit{goto}} \ref{schedule_job} \MyComment{Allow to run anyways}
\State \textbf{\textit{elif}} $J.CPU_{Count} > User_{EntitledCPUCount} - User_{TotalJobsCPUCount}$: \newline \MyComment{Check job's CPU request fits within entitlement}
\State \hspace{0.5cm} $Jobs_{Submitted}.enqueue(J)$ \MyComment{No fit, the job remains in $Jobs_{Submitted}$}
\State \hspace{0.5cm} \textbf{\textit{return}}
\State \textbf{\textit{else}}: \MyComment{User is entitled to run \textit{J}, so make resources available} \newline
\State \hspace{0.5cm} \textbf{\textit{while}} $CPU_{Idle} < J.CPU_{Count}$: \MyComment{Until reaching the entitlements}
\State \hspace{1.0cm} $Job_{checkpoint} \gets Jobs_{Running}.dequeue()$ \MyComment{Checkpoint the least prioritized running jobs}
\State \hspace{1.0cm} \textbf{\textit{if}} $Job_{checkpoint}$ \textit{\textbf{is} checkpointable}: \MyComment{If it is not checkpointable, drop it}
\State \hspace{1.5cm} $Jobs_{Submitted}.enqueue(Job_{checkpoint})$ \MyComment{Checkpointed job back to $Jobs_{Submitted}$}
\State \hspace{1.0cm} $CPU_{Idle} \gets CPU_{Idle} + Job_{checkpoint}.CPU_{Count}$ \MyComment{Update amount of idle CPUs in the system}
\State $Jobs_{Running}.enqueue(J)$ \label{schedule_job} \MyComment{\textit{J} is eligible to run. Schedule it}
\State $CPU_{Idle} \gets CPU_{Idle} - J.CPU_{Count}$ \MyComment{Update amount of idle CPUs in the system} \label{runner_end}
\EndProcedure

\caption{Optimized Memoryless Fair-Share HPC Resource Scheduling using Transparent Checkpoint-Restart Preemption}
\label{algo}
\end{algorithmic}
\end{algorithm*}

In Algorithm \ref{algo} we describe a simplified algorithm, which is based on pre-existing priority queues. They can be governed by any prioritization policy such as FIFO or priority-by-user (lines \ref{jobs_queues_start}--\ref{jobs_queues_end}).
The algorithm starts with system initialization, which occurs only once at startup (lines \ref{sys_init_start}--\ref{sys_init_end}), and with a job initialization procedure which is run every time a new job is initialized (lines \ref{job_init_start}--\ref{job_init_end}).
The \textsc{Memoryless Fair-Share Scheduler} in lines \ref{scheduler_start}--\ref{scheduler_end} iterates over the queue, trying to execute submitted jobs using the main algorithm of \textsc{Memoryless Fair-Share Runner} (lines \ref{runner_start}--\ref{runner_end}). The \textsc{Memoryless Fair-Share Runner} may run the job and may also preempt other jobs, according to the fairness policy described above.

This preemption method lets underutilized entitlements be temporarily used by other entities, and thus improves the utilization over a capping-based system.
Furthermore, an entity can use it to run a single job that is larger than its whole entitlement, without any manual intervention.

Recurrent C/R preemption operations may lead to thrashing.
To resolve this problem we take two approaches: 
First, each job is allowed a \textit{quantum}: A minimal running duration (e.g., 30 minutes). 
The running jobs queue demotes jobs that have been running uninterruptedly for at least a quantum. 
Enlarging the quantum lowers the frequency of C/R operations at the price of non-instantaneous execution of accredited jobs. 
Second, we utilize current developments in distributed file systems~\cite{anderson2019assise, yang2019orion} over non-volatile memory~\cite{izraelevitz2019basic} to reduce the cost of C/R operations: These memory modules are connected to the CPU via the Memory Bus, and can be accessed faster than standard storage.
We intend to use Persistent Memory File Systems~\cite{liu2020survey} (e.g., SplitFS~\cite{splitfs2019}, NOVA-Fortis~\cite{nova2017}, or Assise~\cite{assise2020}) over the Intel Optane™ DataCenter Persistent Memory 
(DCPMM) module~\cite{izraelevitz2019basic}, so that DMTCP will be able to use it transparently. A better approach to resolve this problem is to perform source code modifications to DMTCP, to support Direct Access (DAX) to persistent memory via libraries such as PMDK~\cite{scargall2020introducing}. DAX allows accessing the persistent memory directly from the application. This way, we can utilize the DCPMM to the maximum extent to reduce thrashing cost, as it will serve as a fast non-volatile storage (essential property for checkpointing) and as a memory from which the restart operation could directly start from.  

\section{Related Work}

\subsection{Technical scheduler and transparent C/R}
We follow Rodriguez et al.~\cite{rodriguez2019job} in basing our work on 
SLURM because it is the only scheduler which supports C/R in the form of an API. 
DMTCP is a system-level C/R library which allows to perform C/R operations without any source code modifications; it supports SysV enhancements such as System V shared memory~\cite{dmtcp_faq_2021} which many MPI implementations employ; and it supports InfiniBand~\cite{cao_transparent_2013}. This is in contrast to BLCR~\cite{hargrove2006berkeley}, which requires re-compilation and possibly re-tuning of the kernel module and does not support neither SysV enhancements nor InfiniBand, and CRIU~\cite{criu}, which currently does not even have support for parallel or distributed applications. Furthermore, Rodriguez et al.~\cite{rodriguez2019job} provided two sample scheduling algorithms based on their work. These algorithms do not share the same goal as ours -- which is mainly to increase system utilization -- as one is used to reduce power consumption by scheduling jobs on as least nodes as possible, and the other is used to schedule jobs with higher priority on better hardware.

\subsection{Scheduling methods in SLURM}
Currently, SLURM supports two main scheduling methods: \textit{sched/builtin} which is a simple FCFS algorithm, and \textit{sched/backfill} which is an extension to the basic FCFS algorithm that enables scheduling of jobs that are not at the head of the queue, as long as they do not cause a delay in the scheduling of reserved jobs (jobs which their starting time has already been determined).
Both FCFS and backfill have been studied thoroughly and while being relatively simple, they are used extensively in current supercomputers' job scheduling systems. The main reason is that practical limitations prevent the use of other scheduling algorithms~\cite{Niu2012EmployingCT}. In particular, FCFS-backfill relies on users' estimates for jobs' run-time, which have been proven to be highly inaccurate~\cite{Niu2012EmployingCT, mualem1998, lee2004user}. This evidently leads to very poor efficiency for scheduling systems~\cite{Niu2012EmployingCT}. Niu et al.~\cite{Niu2012EmployingCT} have also shown an example of how checkpointing along with preemptive scheduling can increase the performance of the backfill algorithm, and in this work we follow this approach in practice. To the authors' knowledge, aside from \cite{rodriguez2019job}, this has never been done before.

\subsection{Fairness}
Among others, \textit{fairness} is a common attribute of a job scheduling system in HPC, and it is of utmost importance for HPC users, maybe even more than the productivity of the HPC system~\cite{yuan_pv-easy_2010}.
The concept of \textit{fairness} is well established~\cite{henry_unix_1984}. As \textit{fairness} and performance are in conflict in parallel and distributed job scheduling, previous works either focused on providing fair scheduling or improving performance~\cite{yuan_pv-easy_2010}. A stronger notion of \textit{fairness}, which we did not implement, is \textit{strict fairness}, where no job is delayed by any other job of lower priority.

Our scheduler implements \textit{memoryless fairness}, as opposed to \textit{history-based fair-share}. Schedulers which use this concept re-prioritize jobs according to previous job metrics, with the purpose of reaching a designated level of these metrics (e.g., mean usage or wait time) over a time frame, such that it reflects its entitlements. \textit{History-based fair-share} is less predictable and transparent than \textit{memoryless fair-share}.
As implemented in popular schedulers~\cite{hoopes_slurm_2019, moab_fairshare2019},
\textit{history-based fair-share} is only sensitive to jobs that have been completed and logged by the rolling fair-share calculation.
If a group submits a bunch of jobs quickly, they could potentially fill the entire system with their jobs~\cite{steffen2019better}. Furthermore, since SLURM's implementation of fair-share~\cite{hoopes_slurm_2019, slurm_classic_fairshare} uses a decay factor\footnote{SLURM's fair-share design includes an accounts hierarchy, in which each account is given a normalized relative share of its parent account. For each account, SLURM keeps track of its \textit{actual usage}, and adjusts its \textit{effective usage} accordingly so it represents the account's shares according to the fair-share policy. Both \textit{actual usage}, and \textit{effective usage} are affected by a predefined decay factor, which gives higher weight to a more recent usage of the resources made by each account.}, it is hard to predict when exactly a submitted job will be scheduled.

\section{Conclusion and future work}
In this work we presented the \textit{Optimized Memoryless Fair-Share HPC Resources Scheduling using Transparent Checkpoint-Restart Preemption} (OMFS).
We permanently keep the status quo regarding the \textit{fairness} of the resources distribution -- which is crucial to the users’ work-fashion -- while maximizing the ability of all users to achieve more CPUs and CPU hours for longer periods of time without any non-straightforward costs, penalties or additional human intervention.

In accordance with the notions presented in this paper, we intend to test our scheduler on the NegevHPC~\cite{negevhpc} semi-supercomputer cluster. The first step will be to use SLURM's API using PySlurm~\cite{pyslurm} to build our scheduler, and to evaluate users' behaviour. Next, we plan to build a dedicated SLURM plugin, based on Rodriguez et al.~\cite{rodriguez2019job} work, and to apply our solution to the thrashing problem using non-volatile memory.

\bibliographystyle{plain}
\bibliography{bibliography}

\end{document}